# Protocol Futuring: Speculating Second-Order Dynamics of Protocols in Sociotechnical Infrastructural Futures

### v1

@@ 5.2. From Design Futuring and Infrastructural Futuring to Protocol Futuring
Protocol Futuring is not a substitute for infrastructural inversion, protocol ethnography, or policy prototyping; it complements these approaches by enabling inquiry under conditions where infrastructures are emergent, opaque, or not yet materially instantiated. Infrastructural inversion (bowker1994science) reveals the hidden labor and classificatory schemes underpinning existing systems—but it presupposes a working infrastructure to invert. Protocol ethnography similarly depends on observable practices. Policy prototyping surfaces governance consequences, but is typically constrained to near-future institutional imaginaries. In contrast, Protocol Futuring is most valuable when protocols are nascent, rapidly evolving, or too abstract to have accumulated empirical traces. It provides a systematic way to explore long-range second-order effects when no real-world data yet exists, and to make consequential drift, contestation, or ossification experientially accessible to participants.
@@
Design futuring and critical/speculative design have historically emphasized form-giving: using material artifacts—objects, interfaces, films—to tell stories about possible futures (Dunne2013; fry2009design). Brandt, Binder, and Sanders highlight how materials and design things become shared platforms for participatory exploration and ongoing infrastructuring (Sanders2010framework). In this tradition, speculative artifacts can generate knowledge by embodying tensions, inviting interpretation, and revealing otherwise latent assumptions.
@@
Protocol Futuring diverges from this artifact-centric focus by treating rules, standards, and coordination mechanisms as the primary design material. Instead of asking “What if this object existed?”, we ask “What if this protocol governed interactions?” Yet this does not mean form and material disappear. In our practice, protocols are materialized through a variety of representational forms: pseudo-RFCs, diagrams, decks of rule cards, AI-generated films, and workshop enactments. These are not incidental; they are the forms through which participants grasp, negotiate, and contest protocols. Future work can deepen this connection by explicitly drawing on material speculation—e.g., building physical or mixed-reality artifacts whose behavior is driven by alternative protocol logics—to explore what is gained or lost when rules are embodied in different media. Infrastructural futuring and infrastructural speculations, by contrast, deliberately place lifeworlds, infrastructures, and repair practices at the center of speculative concern (wong2020infrastructural; Harvey2012Topological). These approaches leverage infrastructural inversion to reveal how dependencies, maintenance work, and governance arrangements shape everyday experience. Our method aligns with this orientation but operates at a more meso-level: the scale of protocol regimes that sit between individual artifacts and entire infrastructures. Where infrastructural futuring might imagine future energy grids or mobility systems, Protocol Futuring drills into specific protocol levers within those systems—archival standards, consensus rules, access control schemes—and asks how changes at that level reverberate through infrastructures and lifeworlds over time.
@@ Appendix A Knowledge Futurama Scenario Worksheet
Knowledge Futurama Scenario A
BLUE BUILD ONE – target epoch 2025–2125
Team members (names):
What is the KNOWLEDGE ARTIFACT you want to preserve for 1000 years?
(e.g., genomic data, Bible, Wikipedia…)
What is the TECHNOLOGY STACK for the preservation?
(e.g., IPFS, crystal launched into space, paper copies in vaults, trained LLM on hardened computer…)
Describe all key elements as a bullet list. Try to keep it to 7 items.
What is the SOCIAL MILIEU for the preservation?
(e.g., secret cult of priests, open-source community, universities)
Describe all key elements of the social structure as a bullet list. Try to keep it to 7 items.
@@
RED CRISIS ONE – target event epoch 2120–2125
Team members (names):
Describe the crisis that affects the BLUE BUILD ONE preservation effort in the 5-year period.
Keep it sociologically realistic and technologically plausible. No zombies or time travel. Describe the crisis as a series of dated chronological bullet points spanning the 5 years. The crisis should threaten and reshape BLUE BUILD ONE in unexpected ways, but not destroy it. Keep it to <10<10 bullets.
What is the state of BLUE BUILD ONE at the end of the crisis in 2125?
Describe the state as a set of circumstances/state conditions. Keep the characterization to 7 bullets.
@@
BLUE BUILD TWO – target epoch 2125–2425
Team members (names):
What is the TECHNOLOGY STACK for the preservation?
This can get a little more science fictional than BLUE BUILD ONE, but do not break the laws of physics.
Describe all key elements as a bullet list. Try to keep it to 7 items.
What is the SOCIAL MILIEU for the preservation?
Factor in the milieu described in BLUE BUILD ONE and RED CRISIS ONE. Describe all key elements of the social structure as a bullet list. This can get slightly plausibly transhumanist. Try to keep it to 7 items.
@@
RED CRISIS TWO – target event epoch 2420–2425
Team members (names):
Describe the crisis that affects the BLUE BUILD TWO preservation effort in the 5-year period.
Keep it sociologically realistic and technologically plausible. No zombies or time travel. Describe the crisis as a series of dated chronological bullet points spanning the 5 years. The crisis should threaten and reshape BLUE BUILD TWO in unexpected ways, but not destroy it. Keep it to <10<10 bullets.
What is the state of BLUE BUILD TWO at the end of the crisis in 2625?
Describe the state as a set of circumstances/state conditions. Keep the characterization to 7 bullets.
@@
BLUE BUILD THREE – target epoch 2425–3025
Team members (names):
What is the TECHNOLOGY STACK for the preservation?
This can get even more science fictional than BLUE BUILD TWO, but again do not break the laws of physics.
Describe all key elements as a bullet list. Try to keep it to 7 items.
What is the SOCIAL MILIEU for the preservation?
Factor in the milieu described in BLUE BUILD TWO and RED CRISIS TWO. Describe all key elements of the social structure as a bullet list. This can get even more transhumanist. Try to keep it to 7 items.
@@
RED CRISIS THREE – target event epoch 3020–3025
Team members (names):
Describe the crisis that affects the BLUE BUILD THREE preservation effort in the 5-year period.
Keep it sociologically realistic and technologically plausible. No zombies or time travel. Describe the crisis as a series of dated chronological bullet points spanning the 5 years. The crisis should threaten and reshape BLUE BUILD THREE in unexpected ways, but not destroy it. Keep it to <10<10 bullets.
What is the state of BLUE BUILD THREE at the end of the crisis in 2625?
Describe the state as a set of circumstances/state conditions. Keep the characterization to 7 bullets.
@@ Appendix B Additional Case Studies
B.1. South Beast Asia – Co-Creative Worldbuilding of AI & Protocols in Southeast Asia
@@
Appendix C Disclosure of the usage of LLM
  • Summarize the literature.
  • Polish the existing writing.
  • Proofread for grammar mistakes and spelling.

### v2

@@ 5.2. From Design Futuring and Infrastructural Futuring to Protocol Futuring
Protocol Futuring is not a substitute for infrastructural inversion, protocol ethnography, or policy prototyping; it complements these approaches by enabling inquiry under conditions where infrastructures are emergent, opaque, or not yet materially instantiated. Infrastructural inversion (Bowker, 1994) reveals the hidden labor and classificatory schemes underpinning existing systems—but it presupposes a working infrastructure to invert. Protocol ethnography similarly depends on observable practices. Policy prototyping surfaces governance consequences, but is typically constrained to near-future institutional imaginaries. In contrast, Protocol Futuring is most valuable when protocols are nascent, rapidly evolving, or too abstract to have accumulated empirical traces. It provides a systematic way to explore long-range second-order effects when no real-world data yet exists, and to make consequential drift, contestation, or ossification experientially accessible to participants.
A concrete example from Knowledge Futurama illustrates this complementarity. If one were to study Wikipedia’s current archival practices through infrastructural inversion or protocol ethnography, one might surface the labor of verification, disputes over authority, governance bottlenecks, funding dependencies, and tensions between openness and control. These approaches make visible the hidden work and politics of existing archival infrastructures. What Protocol Futuring uniquely offers is that it extends this analysis by making visible how these same protocol rules transform over centuries of reinterpretation. Through speculative relay and crisis inheritance, participants surfaced dynamics that cannot yet be empirically observed. Rather than revealing what archives are, Protocol Futuring reveals what archive protocols become as they traverse long temporal horizons, shifting social imaginaries, and cumulative reinterpretation.
@@
Design futuring and critical/speculative design have historically emphasized form-giving: using material artifacts—objects, interfaces, films—to tell stories about possible futures (Dunne and Raby, 2013; Fry, 2009). Sanders et al. (2010) highlight how materials and design things become shared platforms for participatory exploration and ongoing infrastructuring. In this tradition, speculative artifacts can generate knowledge by embodying tensions, inviting interpretation, and revealing otherwise latent assumptions.
@@
Protocol Futuring diverges from this artifact-centric focus by treating rules, standards, and coordination mechanisms as the primary design material. Instead of asking “What if this object existed?”, we ask “What if this protocol governed interactions?” Yet this does not mean form and material disappear. In our practice, protocols are materialized through a variety of representational forms: pseudo-RFCs, diagrams, decks of rule cards, AI-generated films, and workshop enactments. These are not incidental; they are the forms through which participants grasp, negotiate, and contest protocols. Future work can deepen this connection by explicitly drawing on material speculation—e.g., building physical or mixed-reality artifacts whose behavior is driven by alternative protocol logics—to explore what is gained or lost when rules are embodied in different media.
Infrastructural futuring and infrastructural speculations, by contrast, deliberately place lifeworlds, infrastructures, and repair practices at the center of speculative concern (Wong et al., 2020; Harvey, 2012). These approaches leverage infrastructural inversion to reveal how dependencies, maintenance work, and governance arrangements shape everyday experience. Our method aligns with this orientation but operates at a more meso-level: the scale of protocol regimes that sit between individual artifacts and entire infrastructures. Where infrastructural futuring might imagine future energy grids or mobility systems, Protocol Futuring drills into specific protocol levers within those systems—archival standards, consensus rules, access control schemes—and asks how changes at that level reverberate through infrastructures and lifeworlds over time.
@@ 6. Conclusion
Acknowledgements.
This work was supported by the Summer of Protocols program, funded by the Ethereum Foundation. We thank Venkatesh Rao, Tim Beiko, and Timber Stinson-Schroff for their guidance and support throughout the program.
Disclosure of the usage of LLM
  • Summarizing literature;
  • Polishing existing writing;
  • Proofreading for grammar and spelling.
@@ Appendix A Knowledge Futurama Scenario Worksheet
BLUE BUILD ONE (2025–2125)
Team members:
What is the KNOWLEDGE ARTIFACT you want to preserve for 1000 years? (e.g., genomic data, Bible, Wikipedia…)
What is the TECHNOLOGY STACK for the preservation? (e.g., IPFS, crystal launched into space, paper copies in vaults, trained LLM on hardened computer…) Describe as a bullet list (7 items max).
What is the SOCIAL MILIEU for the preservation? (e.g., secret cult of priests, open-source community, universities) Describe as a bullet list (7 items max).
@@
RED CRISIS ONE (2120–2125)
Team members:
Describe the crisis affecting BLUE BUILD ONE in this 5-year period. Keep it sociologically realistic and technologically plausible. The crisis should threaten and reshape BLUE BUILD ONE in unexpected ways, but not destroy it (<<10 bullets).
What is the state of BLUE BUILD ONE at the end of the crisis in 2125? Describe as circumstances/state conditions (7 bullets max).
@@
BLUE BUILD TWO (2125–2425)
Team members:
What is the TECHNOLOGY STACK for the preservation? This can get more science fictional than BLUE BUILD ONE, but do not break the laws of physics. Describe as a bullet list (7 items max).
What is the SOCIAL MILIEU for the preservation? Factor in BLUE BUILD ONE and RED CRISIS ONE. Can be slightly transhumanist. Describe as a bullet list (7 items max).
@@
RED CRISIS TWO (2420–2425)
Team members:
Describe the crisis affecting BLUE BUILD TWO in this 5-year period. Keep it sociologically realistic and technologically plausible. The crisis should threaten and reshape BLUE BUILD TWO in unexpected ways, but not destroy it (<<10 bullets).
What is the state of BLUE BUILD TWO at the end of the crisis in 2425? Describe as circumstances/state conditions (7 bullets max).
@@
BLUE BUILD THREE (2425–3025)
Team members:
What is the TECHNOLOGY STACK for the preservation? This can get even more science fictional than BLUE BUILD TWO, but do not break the laws of physics. Describe as a bullet list (7 items max).
What is the SOCIAL MILIEU for the preservation? Factor in BLUE BUILD TWO and RED CRISIS TWO. Can be more transhumanist. Describe as a bullet list (7 items max).
@@
RED CRISIS THREE (3020–3025)
Team members:
Describe the crisis affecting BLUE BUILD THREE in this 5-year period. Keep it sociologically realistic and technologically plausible. The crisis should threaten and reshape BLUE BUILD THREE in unexpected ways, but not destroy it (<<10 bullets).
What is the state of BLUE BUILD THREE at the end of the crisis in 3025? Describe as circumstances/state conditions (7 bullets max).
@@ Appendix B Additional Case Studies
B.1. South Beast Asia